OpenStudy (anonymous):

Solve for x in the proportion 2x/16+30=-1/x x = 8 and x = −8 x = 8 and x = −15 x = −8 x = −3 and x = −5

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

is there an x in the denominator on the left side?

OpenStudy (anonymous):

yea 16 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so is the equation \[\large \frac{2x}{16x+30} = -\frac{1}{x}\]

OpenStudy (anonymous):

the 1 is negitive not the whole thing but the rest is right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok so \[\large \frac{2x}{16x+30} = \frac{-1}{x}\]

jimthompson5910 (jim_thompson5910):

right?

OpenStudy (anonymous):

yea @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\large \frac{2x}{16x+30} = \frac{-1}{x}\] \[\large 2x*x = -1(16x+30)\] \[\large 2x^2 = -16x-30\] \[\large 2x^2+16x+30 = 0\] \[\large 2(x^2+8x+15) = 0\] \[\large x^2+8x+15 = \frac{0}{2}\] \[\large x^2+8x+15 = 0\] \[\large (x+3)(x+5) = 0\] I'll let you finish.

OpenStudy (anonymous):

x = −3 and x = −5? is that right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep, you got it